OpenStudy (anonymous):

one positive number is two more than another. The sum of there squares is 34. Find the smaller number

OpenStudy (anonymous):

one number is x, the other is x +2. x + (x+2)^2 = 34 solve for x

OpenStudy (anonymous):

its actually \[x^{2}+(x+2)^{2}=34\]

OpenStudy (anonymous):

yepp...typo error

OpenStudy (anonymous):

expand (x+2)^2

OpenStudy (anonymous):

2x^2 +4x + 4 = 34

OpenStudy (anonymous):

or x ^2 +2x -15 = 0

OpenStudy (anonymous):

(x +5 )(x - 3) = 0

OpenStudy (anonymous):

\[x^2+(x^2+4x+4)=34\]now combine like terms\[2x^2+4x+4=34\]now move the 34 over and factor to solve for x\[2x^2+4x-30=0\]

OpenStudy (anonymous):

x = -5 or x = 3

OpenStudy (anonymous):

sandy is correct! :)

OpenStudy (anonymous):

thanks!